# The Fragile Nature of Road Transportation Systems

### 01 Feb 2024 v1

@@ 4 Mathematical proof of the fragility of road transportation systems
In this section, we conduct mathematical proof to demonstrate the fragile nature of the transportation systems on both the microscopic and the macroscopic levels. We summarize in Table 1 the structure of this section to prove mathematically the fragile nature of road transportation networks. In the following study, the indicator for the network performance to study the instantaneous disruption onset and between different stable states is the Average Time Spent (ATS), as ATS is the same for each vehicle that entered the network. On the other hand, for the study of the disruption recovery, we use Total Time Spent (TTS), as applied in Zhou and Gayah (2021); Chen et al. (2022); Rodrigues and Azevedo (2019), to better reflect the temporal costs for all the vehicles in this process, as the time spent in the network for different vehicles varies distinctively.
@@ 4.1 Demand disruption
As outlined in Section 2, the presence of a positive second derivative in performance loss concerning the extent of disruption serves as an indication of the transportation system’s fragility. To illustrate the system’s fragility to demand disruption, we analyze the derivatives of time spent relative to the initial disruption demand, either represented by disruption density $k^{\prime}$, or disruption vehicle accumulation $n^{\prime}$. If the system is neither fragile nor antifragile, this measure is expected to exhibit linear growth alongside the disruption. Quoting and in line with the famous statistician George Box, ”All models are wrong, but some are useful.”, we employ both numerical and analytical traffic models to investigate the fragile property under the onset of demand disruption at both microscopic and macroscopic scales.
4.1.1 Onset of demand disruption on the microscopic level
To prove mathematically that transportation systems are fragile on the microscopic level with stable traffic states, we choose the second-degree polynomial FD in Greenshields et al. (1934) as the numerical mathematical model and the two-regime linear FD Daganzo (1994) as the analytical mathematical model.
@@ 4.1.1 Onset of demand disruption on the microscopic level
For the Greenshields second-degree polynomial FD, the following equations describe traffic in a stable state. The traffic flow, density, and speed are denoted as $q$, $k$, and $v$ respectively, while $a$ and $b$ are polynomial coefficients. Then the traffic flow $q$ and average speed $v$ can be determined as in Eq. 2 and Eq. 3, with the speed-density profile shown in Fig. 5(a).
@@
With the sudden onset of a demand disruption with a traffic density of $k^{\prime}$. For a link with a given length of $L$, the ATS and its first and second derivatives over such disruption density $k^{\prime}$ are:
@@
Since the physical meaning of the term $ak^{\prime}+b$ is the average speed of vehicles on this link and should always be a non-negative value while $a$ is negative, thus, both the first and second derivatives of ATS over $k^{\prime}$ are positive, indicating the fragility of transportation systems on a microscopic level.
In this Daganzo two-regime linear model, the speed-density FD can be formulated as the following Eq. 7, as shown in Fig. 5(b).
@@
The second derivative being zero indicates the traffic states before $k_{c}$ are neither fragile nor antifragile. As per Assumption 1, the congested area of the MFD is the study focus for demand disruptions. And when the density is over $k_{c}$, the ATS and its derivatives are:
@@
4.1.2 Onset of demand disruption on the macroscopic level
Likewise, on the macroscopic level, we study the fragile property based on the third-degree polynomial MFD in Geroliminis et al. (2013) and multi-regime linear MFD based on MoC in Daganzo and Geroliminis (2008).
@@ 4.1.2 Onset of demand disruption on the macroscopic level
For MFD approximated with a third-degree polynomial, similar to the Greenshields FD in Eq. 2, we have:
@@
As it should be assumed that Eq. 15 should have real roots, indicating the speed as the root of this second-degree polynomial ought to be a real number, so $b^{2}-4ac>0$ should also hold true. Therefore, the second derivative is always positive.
In MoC, the MFD can be approximated by a series of linear functions. In a similar manner to the Daganzo two-regime linear FD, for any linear function, as in Eq. 13, the second derivative of ATS over density $k$ is defined as:
@@
Therefore, as $c_{i}$ is positive for any cut since the y-intercept should be positive, whether the second derivative is positive or negative depends only on $u_{i}$. Now We group these cuts based on the value of $u_{i}$, and the cuts that intercept the MFD before the critical density $k_{c}$ belong to one group (blue, $u_{i}>0$, as shown in Fig. 6) while the others with intercepts larger than the critical density $k_{c}$ belong to the other group (red, $u_{i}<0$), with an exception in case there’s a cut at the critical density (gray, $u_{i}=0$). Therefore, when the traffic states exceed the critical density, which conforms to Assumption 1, the transportation systems show a fragile property.
@@
4.1.3 Recovery from demand disruption
On the macroscopic level, we study the performance of a road network considering an initial demand as a disruption as well as the capability of this network to absorb and dissipate disruption. Although in real-world cases, a surging traffic demand often lasts over a short period following a certain distribution, for simplicity reasons, we consider this demand as a disruption that happens instantly in the network, which is denoted as $n^{\prime}$ at time $t^{\prime}=0$.
Since the MoC is based on a series of linear functions with decreasing gradients as the vehicle accumulation increases, the complete MoC can be represented into multiple sets of consecutive duo linear functions as Figure 7 shows. The four constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are the slope and y-intercept on the coordinates. While $b_{1}>b_{2}>0$ and are both non-negative values, whether each of $a_{1}$ and $a_{2}$ is positive or negative is yet uncertain. We refer to these two linear functions as the more congested branch and the less congested branch. Here, the critical accumulation $n_{c}$ does not represent the critical point of the entire MFD, but rather the critical accumulation of any consecutive pair of cuts. After a certain period $t_{c}$, the number of vehicles in the network reaches this critical accumulation $n_{c}$. After any period $t>t_{c}$, the vehicle accumulation becomes $n$. We also denote the initial trip completion and critical trip completion as $m_{0}=a_{1}n^{\prime}+b_{1}$ and $m_{c}=a_{1}n_{c}+b_{1}$ respectively.
Figure 7: Simplification of MoC
@@ 4.1.3 Recovery from demand disruption
The duo linear functions representing every two consecutive cuts of the MFD can be formulated into the following Eq. 20, with $a_{2}>a_{1}$ and $b_{1}>b_{2}$:
@@
Therefore, if the traffic states stay only on a specific branch, and when given any initial vehicle accumulation $n_{1}$ at the beginning of a period from $t_{1}$ to $t_{2}$, the number of vehicles $n_{2}$ at the end of this period can be determined through a definite integral:
@@
Here we consider the initial vehicle accumulation after the occurrence of disruption $n^{\prime}$ at the more congested branch and assume after any time $t$, the traffic state is still on the same branch. Then the vehicle accumulation $n$ at any given time $t$ would be:
@@
The TTS can be calculated as the definite integral over this period $t$:
@@
Now we calculate the derivative of TTS over $n^{\prime}$ with $t$ fixed as a constant parameter.
@@
The second derivative of TTS over $n^{\prime}$ is 00, indicating that when the traffic states change only on a single branch, it shows neither fragility nor antifragility.
On the other hand, when the traffic state goes over the critical vehicle accumulation $n_{c}$, meaning it moves from the more congested branch into the less congested branch, the property of fragility or antifragility may be different. As the MFD is a piecewise function, we calculate the TTS separately on both the more and the less congested branches, denoted as $TTS_{1}$ and $TTS_{2}$. Since the critical time $t_{c}$ is still unknown, we need to determine $t_{c}$ first, similar to Eq. 23.
@@
Since TTS is the sum of both $TTS_{1}$ and $TTS_{2}$ each on the more congested and the less congested branches, the second derivative of TTS would also be the sum of the derivatives of $TTS_{1}$ and $TTS_{2}$. The first and second derivatives of $TTS_{1}$ are:
@@
For the vehicle accumulation on the less congested branch from $t_{c}$ to any given time $t$, according to Eq. 24, we have:
@@
According to Assumption 2, we have $m_{0}-q>0$. The second derivative of the whole congestion dissipation process from the more congested branch to the less congested branch would be:
@@
Since $m_{0}-q>0$, and if a transportation system is to be fragile, $\frac{d^{2}TTS}{dn^{\prime 2}}$ should also be positive, and the following equation has to be true:
@@
Since it is assumed that the traffic state transits from the more congested branch into the less congested branch, we have $t>t_{c}$ and $a_{2}>a_{1}$. Therefore, regardless of whether $a_{2}$ is positive or negative, the following relationship always holds:
@@
As the following three terms, $a_{2}-a_{1}$, $(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}$, and $(m_{0}-q)^{\frac{a_{2}}{a_{1}}}$ are all positive, the following relationship always holds true:
@@
And the second derivative of TTS over the disruption vehicle accumulation $n^{\prime}$ is positive. ∎
@@ 4.2 Supply disruption
While a positive second derivative of time spent on traffic demand signifies the fragility of the road transportation network to demand disruptions, establishing a positive second derivative of time spent concerning the magnitude of MFD disruption would demonstrate the fragility from the perspective of supply disruptions. Here we use a supply disruption magnitude coefficient $r$ and the disrupted MFD is expressed as $(1-r)G(n)$. Although real-world MFD may be decreased in various shapes, we use this simple approach as applied in Ambühl et al. (2020) when studying the uncertainties of MFDs. The physical meaning of $(1-r)G(n)$ relates to the decrease of the free-flow speed, for instance, due to snowy weather and icy roads, while the maximal density of the network remains unchanged. The traffic demand at equilibrium before the MFD disruption is $q=G(k_{0})$, or $q=G(n_{0})$. After the supply disruption, as per Assumption 2, the supply disruption magnitude coefficient $r$ is not significantly large so the traffic demand remains below the maximal capacity on the disrupted MFD profile, and the new equilibrium point is $q=(1-r)G(k^{\prime}(r))$, or $q=(1-r)G(n^{\prime}(r))$. It should be noted that, unlike the study of demand disruption, when studying supply disruptions, $k^{\prime}(r)$ is a dependent variable on $r$.
4.2.1 Onset of supply disruption on the microscopic level
First, we study the fragile property of the road transportation networks under the onset of supply disruptions based on the microscopic model Greenshields FD.
@@ 4.2.1 Onset of supply disruption on the microscopic level
The base demand $q$ is constant before and after the onset of the supply disruption.
@@
As $k^{\prime}$ has a physical meaning of density after disruption, hence, it should have real roots with $b^{2}+\frac{4aq}{1-r}$ being positive. And since $1-r$ needs to be positive as well, therefore, $\frac{d^{2}ATS}{dr^{2}}$ is always positive and indicates the fragility property of transportation networks when faced with supply disruption on the microscopic level. Likewise to the proof of demand disruption, the Daganzo FD is a special case of the MoC, so we directly prove the fragility on the macroscopic level with MoC.
4.2.2 Onset of supply disruption on the macroscopic level
Since the proof of fragility under supply disruption with Geroliminis MFD involves solving the roots for cubic equations. For simplicity reasons, we prove only With Daganzo MoC, and we have the traffic demand $q$ and the traffic density of the new equilibrium $k$ as:
@@ 4.2.2 Onset of supply disruption on the macroscopic level
As per Assumption 1, when studying supply disruptions, we focus on the uncongested zone of the MFD, meaning the slope of these relevant cuts is positive so that both the first and second derivatives are positive.
4.2.3 Recovery from supply disruption
To study the possible fragile properties of road transportation networks regarding the recovery process from supply disruptions, we need to combine the conclusions from Sections 4.1.3 and 4.2.2. In Section 4.1.3, we’ve proven the recovery process to be either fragile when the traffic state shifts from the more congested branch to the less congested branch for any two consecutive cuts on the Daganzo MFD, or to be neither fragile nor antifragile when it stays only on one single branch. Therefore, it can be mathematically summarized as $\frac{d^{2}TTS}{dn^{\prime 2}}\geq 0$. Likewise to the proof of positive second derivative, with the assumption of $a_{1}>0$ as the branch is below the critical density following Assumption 1, we can easily prove the first derivative $\frac{dTTS}{dn^{\prime}}$ to be non-negative as well with Eq. 34 and Eq. 40. For the proof of recovery from supply disruptions, as shown in Figure 8, similar to Section 4.2.2, the relationship between the original and new equilibrium points before and after MFD disruptions can be expressed as:
Figure 8: Recovery from supply disruption
@@ 4.2.3 Recovery from supply disruption
Since $u$ and $un_{0}+c$ are both positive, therefore, the first and second derivatives of $n^{\prime}$ over $r$ are positive as well. Additionally, it can be easily proven that when considering the transition from a more congested to a less congested branch, the same conclusion also holds. As $TTS$ is a function of $n^{\prime}$ and $n^{\prime}$ is again a function of $r$, by applying the chain rule, we can get the second derivative of $TTS$ over $r$:
@@
Since all the four components of the Eq. 66 have been proven to be non-negative. Hence $\frac{d^{2}TTS}{dr^{2}}$ is also non-negative and we’ve proven the fragile nature of road transportation systems regarding the recovery process of supply disruptions.

### 04 Mar 2024 v2

@@ 4 Mathematical analysis of the fragility of road transportation systems
In this section, we conduct a mathematical analysis to evaluate the potential fragility of road transportation systems at both microscopic and macroscopic levels. The structure of this section is summarized in Table 1. The reason for not considering the recovery process on the microscopic level is that researchers rarely use FDs to study the recovery from congestion on a link level. And even if we do so, we can also easily get the same conclusion with Proposition 3 and or Proposition 6. In the subsequent study, to investigate the instantaneous disruption onset and between different stable states, the Average Time Spent (ATS) serves as the indicator for the overall performance of a link or a network, since ATS remains constant at any stable traffic state. Conversely, for the examination of disruption recovery, we use Total Time Spent (TTS), as applied in Zhou and Gayah (2021); Chen et al. (2022); Rodrigues and Azevedo (2019), to better reflect the temporal costs for all vehicles in the process, considering that the time spent in the network varies significantly for vehicles entering at different times.
@@ 4.1 Demand disruption
As outlined in Section 3, the presence of a positive second derivative in performance loss concerning the magnitude of disruption serves as an indication of the transportation system’s fragility, therefore, to illustrate the system’s fragility to demand disruption, we analyze the derivatives of time spent, i.e., ATS for the onset of disruptions or TTS for the recovery process, relative to the initial disruption demand, either represented by disruption density $k^{\prime}$ or an initial disruption demand $n^{\prime}$ when we study the relationship between trip completion and vehicle accumulation. If the system is neither fragile nor antifragile, this approach is expected to yield a linearly growing loss in performance alongside the disruption and zero derivatives. Quoting and in line with the famous statistician George Box, “All models are wrong, but some are useful,” we employ both a numerical and an analytical FD/MFD to investigate the fragile properties under the onset of demand disruption at both microscopic and macroscopic scales.
@@
We choose the second-degree polynomial FD in Greenshields et al. (1934) as the numerical traffic model and the two-regime linear FD Daganzo (1994) as the analytical traffic model.
For the Greenshields second-degree polynomial FD, the following equations describe traffic in a stable state with $a$ and $b$ being the polynomial coefficients. The traffic flow $q$ and average speed $v$ can be determined as in Eq. 2 and Eq. 3, with the speed-density profile shown in Fig. 5(a).
@@
With the sudden onset of a demand disruption $k^{\prime}$, for a link with a given length of $L$, the ATS and its first and second derivatives over such disruption density $k^{\prime}$ are:
@@
Since the coefficient $a$ is negative and the physical meaning of the term $ak^{\prime}+b$ is the average speed of vehicles on this link, which should always be positive unless in gridlock, thus, both the first and second derivatives of ATS over $k^{\prime}$ are positive, indicating the fragility of transportation systems on a microscopic level.
In the Daganzo two-regime linear model, the speed-density FD can be formulated as the following Eq. 7, as shown in Fig. 5(b).
@@
With the second derivative being zero, it indicates the traffic states before $k_{c}$ are neither fragile nor antifragile. However, as per Assumption 2, the congested area of the MFD is the study focus for demand disruptions, so now we study the derivatives when $k^{\prime}$ is over $k_{c}$:
@@
Likewise, on the macroscopic level, we study the fragile property based on the third-degree polynomial MFD in Geroliminis et al. (2013) and multi-regime linear MFD based on MoC in Daganzo and Geroliminis (2008). For MFD approximated with a third-degree polynomial, similar to the Greenshields FD in Eq. 2, we have:
@@
The average speed has to be a real number, indicating Eq. 15 should have real roots, so $b^{2}-4ac>0$ should hold true. Therefore, the derivatives are positive.
The MFD derived with MoC can be approximated by a series of linear functions. Likewise to the Daganzo two-regime linear FD as in Eq. 13, for any linear function, the second derivative is:
@@
As coefficient $c_{i}$ is positive for any cut since the y-intercept should always be positive by definition of MoC, whether the second derivative is positive or negative depends solely on $u_{i}$. The cuts that intercept the MFD before the critical density $k_{c}$ exhibit antifragile properties ($u_{i}>0$, in blue as shown in Fig. 6) while the others with intercepts larger than the critical density $k_{c}$ show fragile responses ($u_{i}<0$, in red), with an exception in case there’s a cut at the critical density ($u_{i}=0$, in gray). Conforming to Assumption 2, for demand disruptions, we focus on the cuts with intercepts larger than the critical density $k_{c}$ ($u_{i}<0$). The second derivative for these cuts is positive indicating a fragile property.
@@
According to Assumption 1, we simplify this surging demand as a disruption that takes place instantly in the network, denoted as $n^{\prime}$ at time $t^{\prime}=0$. As the MoC is composed of a series of linear functions with decreasing gradients as the vehicle accumulation increases, the multi-linear regimes can be represented into multiple sets of consecutive duo linear functions as Fig. 7 shows. The four constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are the slope and y-intercept on the coordinates for the two cuts, with $a_{2}>a_{1}$ and $b_{1}>b_{2}>0$. We refer to these two linear functions as the more congested branch and the less congested branch. Also, the critical accumulation $n_{c}$ here does not represent the critical point of the entire MFD, but rather the critical accumulation of any two consecutive cuts. After a certain period $t_{c}$, the number of vehicles in the network reaches this critical accumulation $n_{c}$. And after any period $t>t_{c}$, the vehicle accumulation becomes $n$. We also denote the initial trip completion and critical trip completion as $m_{0}=a_{1}n^{\prime}+b_{1}$ and $m_{c}=a_{1}n_{c}+b_{1}=a_{2}n_{c}+b_{2}$ respectively.
Figure 7: Simplification of MoC
Any two consecutive cuts of the MFD can be formulated into the following Eq. 20:
@@
When the traffic states move only along a single branch, and given any initial vehicle accumulation $n_{1}$ at the beginning of a period from $t_{1}$ to $t_{2}$, the number of vehicles $n_{2}$ at the end of this period can be determined as:
@@
Therefore, with the disruption accumulation $n^{\prime}$, and when the traffic states are on the same branch. After any time $t$, the vehicle accumulation $n$ would be:
@@
The TTS in this case can be calculated as:
@@
Now we calculate the derivatives of TTS considering $t$ as any positive constant.
@@
The second derivative of TTS is 00, indicating that when the traffic states move only along a single branch, it shows neither fragility nor antifragility.
On the other hand, when the traffic state goes over the critical vehicle accumulation $n_{c}$, and since the MoC is a piecewise function, we calculate the TTS separately on both the more congested and the less congested branches, denoted as $TTS_{1}$ and $TTS_{2}$. Since the critical time $t_{c}$ is still unknown, we need to determine $t_{c}$ first, similar to Eq. 23.
@@
Since TTS is the sum of $TTS_{1}$ and $TTS_{2}$, the second derivative of TTS would also be the sum of the derivatives. The derivatives for $TTS_{1}$ are:
@@
According to Eq. 24, the vehicle accumulation on the less congested branch from $t_{c}$ to $t$ would be:
@@
The second derivative of the whole process can be written as:
@@
As per Assumption 3, we have $m_{0}-q>0$, so if a transportation system is to be fragile, $d^{2}TTS/dn^{\prime 2}$ should also be positive, and the following equation has to be true:
@@
Since $t>t_{c}$ and $a_{2}>a_{1}$, regardless of whether $a_{2}$ is positive or negative, the following relationship always holds:
@@
As the following three terms, $a_{2}-a_{1}$, $(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}$, and $(m_{0}-q)^{\frac{a_{2}}{a_{1}}}$ are all positive, the following relationship is true:
@@
The second derivative of TTS over the disruption vehicle accumulation $n^{\prime}$ is positive, which indicates the fragility. ∎
@@ 4.2 Supply disruption
While a positive second derivative of ATS or TTS on traffic demand signifies the fragility of the road transportation network to demand disruptions, establishing a positive second derivative of time spent concerning the magnitude of FD/MFD disruption would demonstrate the fragility from the perspective of supply disruptions. Here we use a supply disruption magnitude coefficient $r$ and the disrupted FD/MFD is expressed as $(1-r)G(n)$. Although real-world MFD may be decreased in various shapes, we use this simple approach as applied in Ambühl et al. (2020) when studying the uncertainties of MFDs. The physical meaning of $(1-r)G(n)$ relates to the decrease of the free-flow speed due to, e.g., snowy weather and icy roads, with the maximal density of the network remaining unchanged. The traffic demand at equilibrium before the MFD disruption is $q=G(k_{0})$, or $q=G(n_{0})$. After the supply disruption, as per Assumption 3, the supply disruption magnitude coefficient $r\in[0,1)$ is not significantly large so the traffic demand remains below the maximal capacity on the disrupted MFD profile, and the new equilibrium point is $q=(1-r)G(k^{\prime}(r))$, or $q=(1-r)G(n^{\prime}(r))$. It should be noted that, unlike the study of demand disruption, when studying supply disruptions, $k^{\prime}(r)$ is a dependent variable on $r$.
@@
First, we study the fragile properties of the road transportation systems under the onset of supply disruptions based on the microscopic model Greenshields FD. Here, the base demand $q$ is constant before and after the onset of the supply disruption, and can be used to build the relationship between the two stable states before and after the supply disruption.
@@
As $k^{\prime}$ has a physical meaning of disruption density, so it should have real roots with $b^{2}+4aq/1-r$ being positive. And since $01-r>$ 0, therefore, $d^{2}ATS/dr^{2}$ is always positive and indicates the fragility property of transportation systems when faced with supply disruption on the microscopic level. Likewise to the proof of demand disruption, the Daganzo FD is a special case of the MoC, so we directly prove the fragility on the macroscopic level with MoC illustrated in the following section.
@@
As the proof of fragility under supply disruption with cubic polynomial MFDs involves solving the roots for cubic equations, for simplicity reasons, we prove only with Daganzo MoC. Since the traffic demand $q$ is constant, the traffic density of the new stable state $k^{\prime}(r)$ would be:
@@
As per Assumption 2, when studying supply disruptions, we focus on the uncongested zone of the MFD, meaning the slope of these relevant cuts is positive so that both derivatives are positive.
@@
To study the possible fragile properties of road transportation networks regarding the recovery process from supply disruptions, we need to combine the conclusions from Proposition 3 and Proposition 5. In Proposition 3, we’ve proven the recovery process to be fragile when the traffic state shifts from the more congested branch to the less congested branch for any two consecutive cuts on the Daganzo MFD, or to be neither fragile nor antifragile when it stays only on one single branch. Therefore, it can be mathematically summarized as $d^{2}TTS/dn^{\prime 2}\geq 0$. Following Assumption 2, $a_{1}>0$ holds as the branch is below the critical density, we can easily prove the first derivative $dTTS/dn^{\prime}$ to be non-negative as well with Eq. 34 and Eq. 40 through the same procedure as proving the second derivative to be positive. And for the proof of recovery from supply disruptions, as shown in Fig. 8, likewise to Proposition 5, the relationship between the original and new equilibrium points before and after MFD disruptions can be expressed as:
Figure 8: Recovery from supply disruption
@@
Since $u$ and $un_{0}+c$ are both positive, the derivatives of $n^{\prime}$ over $r$ are positive as well. Additionally, it can be easily proven that when considering the transition from a less congested to a more congested branch, the same conclusion also holds. As $TTS$ is a function of $n^{\prime}$ and $n^{\prime}$ is again a function of $r$, by applying the chain rule, we can get the second derivative of $TTS$ over $r$ as:
@@
Because all the four components of the Eq. 66 have been demonstrated above to be non-negative, thus $d^{2}TTS/dr^{2}$ is also non-negative and we’ve proven the fragile nature of road transportation systems regarding the recovery process of supply disruptions.